Terrain Rendering for a Geographical Information System

Paul Coddington

Northeast Parallel Architectures Center

November 1994

1 Terrain Rendering

1.1 Rendering of Terrain

Terrain rendering is a special case of 3-D volume rendering:

• The objects are static – only the viewer moves.

• There is only a single simply connected object – the land surface.

• A terrain model is basically a 2-D surface with a variable elevation in the third dimension, which is much simpler than a general 3-D problem. Terrain rendering is often called a 2.5 dimensional problem.

For these reasons, there are a number of optimizations and simplifications that can be made for a terrain rendering program that cannot be done for a general purpose volume rendering program.

1.2 Portability and Scalable Rendering

For best results, the rendering must be done on a high-performance computer. However the renderer should be portable, so it will run on parallel computers with different numbers of nodes, high-end graphics workstations, and even low-end platforms like Macs.

Our strategy is to make a scalable algorithm that will always run in near real time, by (gracefully) sacrificing image quality when there is less compute power available. This can be done relatively easily using polygon-based rendering by coarsening, i.e. reducing the number of polygons used.

This method requires handling two key problem areas which are interesting research topics:

• Keeping a constant frame rate for different computers and different scenes